luhivqqcherry:

help with mathh !! . Write the nth term formula: 5 + 11 + 17 + 23 + … + 35 a) t(n) = 5 + 6(n – 1) b) t(n) = 5(6)n-1 c) t(n) = -1 + 6(n - 1) d) t(n) = 5 + 6n

Midnight97:

Let's see what I can do

ShadowKid3:

The sequence forms an arithmetic sequence, displaying a common difference of 6 between consecutive terms. To derive the nth term formula, we employ the formula for the nth term of an arithmetic sequence, which is t(n) = a + (n-1)d. Here, t(n) represents the nth term, a signifies the first term, n denotes the position of the term, and d stands for the common difference. In this instance, the first term (a) equals 5, while the common difference (d) equals 6, resulting in the nth term formula t(n) = 5 + 6(n - 1). So, the right answer is a) t(n) = 5 + 6(n – 1).

luhivqqcherry:

@shadowkid3 wrote:
The sequence forms an arithmetic sequence, displaying a common difference of 6 between consecutive terms. To derive the nth term formula, we employ the formula for the nth term of an arithmetic sequence, which is t(n) = a + (n-1)d. Here, t(n) represents the nth term, a signifies the first term, n denotes the position of the term, and d stands for the common difference. In this instance, the first term (a) equals 5, while the common difference (d) equals 6, resulting in the nth term formula t(n) = 5 + 6(n - 1). So, the right answer is a) t(n) = 5 + 6(n – 1).
ahhh tytyytysmm !! .